# Feature Representation Transferring to Lightweight Models via Perception Coherence

### v1

@@ 4.2 Learning Metric Evaluation on Neural Networks
In this section, we perform experiments on CIFAR10 cifar10 and the subset of the first 30 classes of CUB-200 wah2011caltech with very small student models. Following the previous works (e.g. passalis2018learning ; passalis2020probabilistic ; passalis2020heterogeneous ), we evaluate the quality of the learned representation using the content-based image retrieval setup. We used the official code released with the work passalis2020heterogeneous to evaluate different metrics, including the interpolated mean Average Precision (mAP) (at the standard 11-recall points) and the top-$k$ precision. For a fair comparison, we use the same teacher and student model architecture as in passalis2020heterogeneous . The student models are very small, composed of only 3 convolutional layers and a fully connected layer. All details are in Appendix E.2. To conduct the experiments, for each dataset CIFAR10 and CUB-200, we first train the teacher model on the training set. Then, the transfer set is only the training set without label. Then, the retrieval setup is used to evaluate the quality of the learned representation. The database contains the training set and the test set is used for evaluating the performance of each model. If the student is trained to learn useful features, then the test examples should be close to the training examples (database) of the same class. Hence, a higher retrieval score indicates a better quality of the learned representation. The results for CIFAR10 and CUB-200 are shown in Tables 4.1 and 4.2, respectively. Our method is also compared with classical methods, including the standard knowledge distilling (KD) hinton2015distilling , FitNet romero2014fitnets , metric knowledge transfer (MKT) yu2019learning , probabilistic knowledge transfer (PKT) passalis2018learning and Heterogeneous Knowledge Distillation (HKD) passalis2020heterogeneous . The results displayed for all these methods in Tables 4.1 and 4.2 are extracted from passalis2020heterogeneous . Therein, one used the same teacher and student models. For FitNet, MKT, PKT, there are two versions: the first one applies only on a single penultimate layer, while the second one applies on multiple layers (denoted by -H).
Table 4.1: Learning Metric Evaluation on CIFAR10.
@@
KD hinton2015distilling No 40.53 58.56
FitNet romero2014fitnets No 48.99 62.42
MKT yu2019learning No 38.20 52.72
PKT passalis2018learning No 51.56 62.50
@@
HKD passalis2020heterogeneous Yes 53.06 64.24
Table 4.2: Learning Metric Evaluation on a subset of the first 30 classes of CUB200.
@@
KD hinton2015distilling No 18.55 26.57
FitNet romero2014fitnets No 15.98 23.41
MKT yu2019learning No 13.39 20.59
PKT passalis2018learning No 18.57 26.70
@@
HKD passalis2020heterogeneous Yes 19.01 27.67
From Tables 4.1 and 4.2, we observe that MKT method provides the worst results. This method is composed of 2 main ingredients: (1) minimizing the distance between the student and teacher features and (2) enforcing the student model to result in similar distances between the data points as the teacher model (in the feature space). Hence, in our setting of very small student models, this brute force approach provides the worst results. Our method also outperforms by a large gap the standard approaches of KD (using the soft label) or FitNet (using the distance between the student and teacher features). More advanced methods such as PKT (based on a kernel method) performs better, but it is still outperformed by our method. Besides, HKD is a more advanced version of PKT, where one uses a particular mechanism to transfer the features. However, our method still provides better results. Notice that our method is applied only on the penultimate layer, but outperforms the other methods using multiple layers. Especially on CUB-200, our method outperforms the others by a wide margin. This proves the effectiveness of our method for feature representation transferring.
@@
In the above section, we show how our method helps the student model to produce useful features. In this section, we demonstrate that by transferring the feature representation, our method boosts significantly the performance of the student model in the context of the classification task on CIFAR100 krizhevsky2009learning . While different methods have been proposed and tailored specifically for the classification, we focus mainly on state-of-the-art feature representation transferring methods. Notice that different methods (such as FitNet romero2014fitnets ) apply the transferring on different intermediate layers. However, we propose to apply our method only on the penultimate layer and on the logit one (before applying the softmax function). We also fix all the hyperparameters for training, without further tuning to show the simplicity of our method. All details can be found in Appendix E.3. While different methods have been proposed to effectively transfer knowledge between model of the same architecture type (e.g. ResNet50 $\to$ ResNet18), we intentionally choose pairs of models having different architecture types. More precisely, we use 3 pairs of teacher and student models: ResNet-50 $\to$ MobileNetV2 and ResNet-32x4 $\to$ ShuffleNetV1 / ShuffleNetV2. We follow the same setup and use the same pretrained teacher models officially released with the work tiancontrastive for fair comparisons.
[…]
@@
KD hinton2015distilling 67.35$\pm$0.32 74.07$\pm$0.19 74.45$\pm$0.27
FitNet romero2014fitnets 63.16$\pm$0.47 73.59$\pm$0.15 73.54$\pm$0.22
VID ahn2019variational 67.57$\pm$0.28 73.38$\pm$0.09 73.40$\pm$0.17
RKD park2019relational 64.43$\pm$0.42 72.28$\pm$0.39 73.21$\pm$0.28
PKT passalis2018learning 66.52$\pm$0.33 74.10$\pm$0.25 74.69$\pm$0.34
CRD tiancontrastive 69.11$\pm$0.36 75.11$\pm$0.32 75.65$\pm$0.10
Ours 68.64$\pm$0.10 75.62$\pm$0.28 76.04$\pm$0.25
@@
From Table 4.3, we see that our method boosts significantly the performance of the student models (when not using any KD technique). It also provides much better results than standard techniques such as KD or FitNet. This proves its effectiveness for extracting knowledge from the teacher model. Moreover, it outperforms all the other methods on ResNet-32x4/ShuffleNetV1 and ResNet-32x4 /ShuffleNetV2 benchmarks. On the experiment ResNet-50/ MobileNetV2, our method is slightly outperformed by the method CRD, but outperforms all the other methods. Besides, our method systematically outperforms the PKT method based on kernel techniques (see Section 2). This reinforces our intuition that our method is more effective than applying the same kernel for all points, as discussed in Section 2. Other methods based on mutual information such as Variational Information Distillation (VID) is also outperformed by ours. The method of Contrastive Representation Distillation (CRD) is also based on an information-theoretic framework. Our method provides better or on-par performance compared to CRD. All these remarks prove the effectiveness of our method for KD.
@@
Finally, notice that some different knowledge distilling methods are tailored specifically for classification (e.g. huang2022knowledge ). However, our main objective is to study methods for feature representation transferring, as this class of methods is more generic and applicable on a wider range of tasks (e.g. retrieval). Through our experiments, we aim to show that a good feature representation transfer allows us to effectively distill the knowledge from a neural network.
@@ 5 Discussion and Conclusion
Preserving dissimilarity ranks is a flexible requirement, as it does not enforce a full copy of the geometry. However, a question remains theoretically unresolved in this work: How well can a given student model preserve these rankings (as a function of its capacity)? This question remains open in the present work and will be explored in future theoretical works. Besides, notice that by considering the dissimilarity ranking, our method provides structural information of the data manifold. This is related to topological data analysis (TDA), which aims to understand the global data structure. The intersection of TDA and Deep Learning is referred to as topological deep learning (TDL) zia2023topological . Topology studies invariant properties under continuous deformation. In our case, as long as the ranking of relatives distances remains unchanged, we also have invariance under local and global deformation. This allows a better flexibility and stability. One main difference of our method is that, topology studies a configuration based on deterministic approaches, whereas our method proposes a more relaxed view by introducing a probabilistic framework. Besides, by integrating neural networks into this framework, our work is the first attempt to develop TDL for representation transfer and KD.

### v2

@@ 5.2 Learning Metric Evaluation on Neural Networks
In this section, we perform experiments on CIFAR10 Krizhevsky et al. (2009) and the subset of the first 30 classes of CUB-200 Wah et al. (2011) with very small student models. Following the previous works (e.g. Passalis & Tefas (2018a); Passalis et al. (2020b; a)), we evaluate the quality of the learned representation using the content-based image retrieval setup. We used the official code released with the work Passalis et al. (2020a) to evaluate different metrics, including the interpolated mean Average Precision (mAP) (at the standard 11-recall points) and the top-$k$ precision (Christopher et al., 2008). For a fair comparison, we use the same teacher and student model architecture as in Passalis et al. (2020a). The student models are very small, composed of only 3 convolutional layers and a fully connected layer. All details are in Appendix E.2. To conduct the experiments, for each dataset CIFAR10 and CUB-200, we first train the teacher model on the training set. Then, the transfer set is only the training set without label. Then, the retrieval setup is used to evaluate the quality of the learned representation. The database contains the training set and the test set is used for evaluating the performance of each model. If the student is trained to learn useful features, then the test examples should be close to the training examples (database) of the same class. Hence, a higher retrieval score indicates a better quality of the learned representation. The results for CIFAR10 and CUB-200 are shown in Tables 1 and 2, respectively. Our method is also compared with classical methods, including the standard knowledge distilling (KD) Hinton et al. (2015), FitNet Romero et al. (2014), metric knowledge transfer (MKT) Yu et al. (2019), probabilistic knowledge transfer (PKT) Passalis & Tefas (2018a) and Heterogeneous Knowledge Distillation (HKD) Passalis et al. (2020a). The results displayed for all these methods in Tables 1 and 2 are extracted from Passalis et al. (2020a). Therein, one used the same teacher and student models. For FitNet, MKT, PKT, there are two versions: the first one applies only on a single penultimate layer, while the second one applies on multiple layers (denoted by -H).
Table 1: Learning Metric Evaluation on CIFAR10.
@@
KD (Hinton et al., 2015) No 40.53 58.56
FitNet (Romero et al., 2014) No 48.99 62.42
MKT (Yu et al., 2019) No 38.20 52.72
PKT (Passalis & Tefas, 2018a) No 51.56 62.50
@@
HKD (Passalis et al., 2020a) Yes 53.06 64.24
Table 2: Learning Metric Evaluation on a subset of the first 30 classes of CUB200.
@@
KD (Hinton et al., 2015) No 18.55 26.57
FitNet (Romero et al., 2014) No 15.98 23.41
MKT (Yu et al., 2019) No 13.39 20.59
PKT (Passalis & Tefas, 2018a) No 18.57 26.70
@@
HKD (Passalis et al., 2020a) Yes 19.01 27.67
From Tables 1 and 2, we observe that MKT method provides the worst results. This method is composed of 2 main ingredients: (1) minimizing the distance between the student and teacher features and (2) enforcing the student model to result in similar distances between the data points as the teacher model (in the feature space). Hence, in our setting of very small student models, this brute force approach provides the worst results. Our method also outperforms by a large gap the standard approaches of KD (using the soft label) or FitNet (using the distance between the student and teacher features). More advanced methods such as PKT (based on a kernel method) performs better, but it is still outperformed by our method. Besides, HKD is a more advanced version of PKT, where one uses a particular mechanism to transfer the features. However, our method still provides better results. Notice that our method is applied only on the penultimate layer, but outperforms the other methods using multiple layers. Especially on CUB-200, our method outperforms the others by a wide margin. This proves the effectiveness of our method for feature representation transferring.
@@
In the above section, we show how our method helps the student model to produce useful features. In this section, we demonstrate that by transferring the feature representation, our method boosts significantly the performance of the student model in the context of the classification task on CIFAR100 Krizhevsky (2009). While different methods have been proposed and tailored specifically for the classification, we focus mainly on state-of-the-art feature representation transferring methods. Notice that different methods (such as FitNet Romero et al. (2014)) apply the transferring on different intermediate layers. However, we propose to apply our method only on the penultimate layer and on the logit one (before applying the softmax function). We also fix all the hyperparameters for training, without further tuning to show the simplicity of our method. All details can be found in Appendix E.3. While different methods have been proposed to effectively transfer knowledge between model of the same architecture type (e.g. ResNet50 $\to$ ResNet18), we intentionally choose pairs of models having different architecture types. More precisely, we use 3 pairs of teacher and student models: ResNet-50 $\to$ MobileNetV2 and ResNet-32x4 $\to$ ShuffleNetV1 / ShuffleNetV2. We follow the same setup and use the same pretrained teacher models officially released with the work Tian et al. (2020) for fair comparisons.
Table 3: Test classification accuracy ($\%$) on CIFAR100 using different knowledge distillation methods. Results of other methods are extracted from Huang et al. (2022), where one used the same setups.
@@
[…]
@@
From Table 3, we see that our method boosts significantly the performance of the student models (when not using any KD technique). It also provides much better results than standard techniques such as KD or FitNet. This proves its effectiveness for extracting knowledge from the teacher model. Moreover, it outperforms all the other methods on ResNet-32x4/ShuffleNetV1 and ResNet-32x4 /ShuffleNetV2 benchmarks. On the experiment ResNet-50/ MobileNetV2, our method is slightly outperformed by the method CRD, but outperforms all the other methods. Besides, our method systematically outperforms the PKT method based on kernel techniques (see Section 2). This reinforces our intuition that our method is more effective than applying the same kernel for all points, as discussed in Section 2. Other methods based on mutual information such as Variational Information Distillation (VID) is also outperformed by ours. The method of Contrastive Representation Distillation (CRD) is also based on an information-theoretic framework. Our method provides better or on-par performance compared to CRD. All these remarks prove the effectiveness of our method for KD.
@@
Finally, notice that some different knowledge distilling methods are tailored specifically for classification (e.g. Huang et al. (2022)). However, our main objective is to study methods for feature representation transferring, as this class of methods is more generic and applicable on a wider range of tasks (e.g. retrieval). Through our experiments, we aim to show that a good feature representation transfer allows us to effectively distill the knowledge from a neural network.
5.4 Ablation study: effect of mini-batch size
In Section 4.1, we established a theoretical result showing that the global perception coherence level (GPCL), when estimated using mini-batches of size $B$, converges in expectation to its true value at a rate of $O(1/\sqrt{B})$ (Theorem 4.1). We now turn to an empirical validation of this result. To this end, we employ the estimator defined in Definition 4.1 with varying batch sizes, and evaluate it on trained student models from the CIFAR10 and CUB200 experiments. The outcomes are summarized in Figure 6 and Table 4.
From Table 4, we observe that for very small batch sizes (e.g., $B=4$), the estimated GPCL deviates significantly from the reference value, reflecting the limited information available in such small batches. As the batch size increases (e.g., $B=16$ or $B=64$), the estimated GPCL rapidly converges toward the true value. In fact, with $B=32$ or $B=64$, the estimator already provides a stable and reliable approximation. For instance, in the CIFAR10 case with 10,000 test examples, using $B=64$ yields an accurate estimate with low variance, thereby confirming the predicted convergence rate with respect to batch size.
Table 4: Estimation of global perception coherence level, using different mini-batch sizes.
Mini-batch size 4 8 16 32 64 128 256 full dataset
CIFAR10 $0.8943\pm 0.00098$ $0.8377\pm 0.00061$ $0.8131\pm 0.00032$ $0.8021\pm 0.00029$ $0.7969\pm 0.00028$ $0.7944\pm 0.00017$ $0.7931\pm 0.00014$ 0.7919
CUB200 subset $0.8919\pm 0.00271$ $0.8312\pm 0.00114$ $0.8073\pm 0.00085$ $0.7959\pm 0.00136$ $0.7908\pm 0.00081$ $0.7878\pm 0.00059$ $0.7874\pm 0.00057$ 0.7859
Figure 6: Estimated global perception coherence level at different batch sizes.
To further investigate the qualitative impact of batch size, we conduct experiments on toy datasets where representations are transferred from $\mathbb{R}^{3}\to\mathbb{R}^{2}$ (Figure 7) and from $\mathbb{R}^{2}\to\mathbb{R}^{2}$ (Figure 8). Here, we explicitly vary $B$ and analyze the resulting student configurations. When $B$ is too small (e.g., $B=3$), the learned configuration fails to preserve the structural coherence of the reference configuration. This is expected, as such small batches provide only very limited relational information between points. Starting from $B=7$, however, the learned configurations begin to recover the global structure, and from $B=15$ onward, the results become stable.
These findings suggest that excessively large batch sizes are not necessary for capturing global coherence. Indeed, during training, each epoch involves drawing random mini-batches of size $B$, and within each batch, every reference point is compared to B1B-1 others. Due to random sampling, across successive epochs each point is paired with different comparison sets. As a result, the model gradually aggregates sufficient global structural information, even when $B$ is relatively small compared to the dataset size. This explains why reliable approximations of GPCL and structurally coherent student configurations can be obtained without resorting to very large batch sizes.
Figure 7: Ablation study of the effect of batch size on $3D\to 2D$ Gaussian dataset, where we train with different batch sizes. Plots from left to right in each row represent evolution of the learned configuration during training.
Figure 8: Ablation study of the effect of batch size on two-moon dataset, where we train with different batch sizes.
5.5 Ablation study: Student model size
In this section, we investigate the influence of student model size on transfer performance. Recall that in the CIFAR10 and CUB200 experiments, the baseline student architecture consists of three convolutional layers (conv1/conv2/conv3), each followed by a ReLU non-linearity, and a final fully connected layer. To explicitly examine the effect of model size, we construct progressively smaller student variants by truncating the network at different depths and applying Global Average Pooling (GAP) [^2] Formally, GAP: $\mathbb{R}^{H\times W\times C}\mapsto\mathbb{R}^{C}$, such that for $z\in\mathbb{R}^{H\times W\times C}$, for $k\in\{1,2,\cdots,C\}$, $\text{GAP}(z)_{k}=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}z_{ijk}$. over the two spatial dimensions to obtain the final representation:
  • Model 1 (one layer): conv1 + GAP.
  • Model 2 (two layers): conv1 + conv2 + GAP.
  • Model 3 (three layers): conv1 + conv2 + conv3 + GAP.
  • Model 4: the full student model (original architecture).
The results are reported in Table 5. For both datasets, we observe a consistent trend: as the student model becomes larger, the global perception coherence level (GPCL) increases. This indicates that deeper models are more capable of preserving the ranking structure induced by the teacher’s feature space.
Moreover, we note a strong correlation between GPCL and downstream performance metrics such as mean Average Precision (mAP) and top-$k$ accuracy. In particular, models with higher GPCL systematically achieve better classification performance. This empirical evidence reinforces the role of GPCL as a meaningful and effective objective for guiding teacher–student learning. It also highlights a practical limitation: when the student is excessively small, its limited representational capacity hampers the preservation of structural coherence, leading to degraded performance.
Table 5: Performance of different models, with different sizes, with the same teacher model.
Data set Metric One layer Two layers Three Layers Full model
CIFAR10 Global PC Level (on test set) 0.7015 0.7491 0.7700 0.7919
mAP 17.08 26.52 37.69 54.25
top-100 20.43 35.36 50.29 65.00
CUB200 subset Global PC Level (on test set) 0.7015 0.7266 0.7602 0.7859
mAP 6.63 8.66 18.88 28.42
top-10 10.35 13.38 27.28 36.55
@@ 6 Discussion and Conclusion
Preserving dissimilarity ranks constitutes a flexible requirement, as it does not enforce a complete replication of the teacher’s geometry. Nevertheless, an important theoretical question remains open: to what extent can a given student model preserve these rankings as a function of its representational capacity? Addressing this question lies beyond the scope of the present work but represents a promising direction for future theoretical investigations.
Why rankings? Generally, in the teacher model’s feature space, smaller distances (or dissimilarities) often reflect higher semantic similarity. By teaching the student model to preserve the relative rankings, it learns to order semantic dissimilarities in a way that aligns with how the teacher model perceives the input semantic dissimilarity. In order to replicate this ordering, the student model has to understand features that capture similar semantic content—thus encouraging the learning of meaningful and transferable representations.
By considering dissimilarity rankings, our method implicitly captures structural information about the underlying data manifold. This perspective can be viewed as a form of topology-aware representation transfer. Topology is concerned with invariant properties of a configuration under continuous deformations. Analogously, in our framework, as long as the relative distance orderings are preserved, the learned representation remains invariant under local or global distortions. This provides additional flexibility and stability, distinguishing our approach from geometry-preserving methods. A notable difference from classical topological approaches lies in the probabilistic framework that we introduce. While topology typically characterizes data configurations deterministically, our method relaxes this view by embedding ranking preservation in a probabilistic setting. Furthermore, by integrating neural networks into this framework, we open the path toward topology-aware knowledge transfer in deep learning.